OpenStudy (anonymous):

solve: x+y=13 4x+2y=13 by substitution or elimination technique

OpenStudy (anonymous):

hey can you please check this problem In a farmyard there are only chickens and sheep. There are 13 animals altogether and between them they have 34 legs. Use simultaneous equations to determine how many chickens and sheep there are. same problem not sure if equation are correct

hero (hero):

x + y = 13 2x + 4y = 34

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how to solve them

hero (hero):

Same way I solved the other one

OpenStudy (anonymous):

but you -7

hero (hero):

You set up that first one wrong then

hero (hero):

Erasing...one minute please

hero (hero):

x + y = 13 2x + 4y = 34 y = 13 - x y = (34 - 2x)/4 y = y 4(13 - x) = 34 - 2x 52 - 4x = 34 - 2x 2x = 18 x = 18/2 x = 9 y = 4

hero (hero):

x = chicken y = sheep

OpenStudy (anonymous):

so there are chicken and 4 sheep. how do I prove that their legt

OpenStudy (anonymous):

how do I prove that their legs are correct

hero (hero):

2(9) + 4(4) = 34 18 + 16 = 34 34 = 34

OpenStudy (anonymous):

how did

OpenStudy (anonymous):

how did you get the number of sheep(4)